The diagonal of a rectengular field is 60 more than the shorter side if longer side is 30 more than the shorter side find the sides of the field

Grade:10

3 Answers

Aman Kashyap
6 years ago
By prepainga diagram, sides x, x+30 and x+60 form a right angle triange inside the rectangle. We can apply pythagoras theorem to this triangle and obtain the value of x. (x^2) + (x+30)^2 = (x+60)^2. By solving this equations and cancelling some terms we are left with a simple quadratic equation. Which is x^2 -60x -2700=0. This can be factorized be splitting the mid term and the value of x turns out to be 90. Hence shortest side =90.
sachin
6 years ago
sides x, x+30 and x+60 form a right angle triange inside the rectangle. We can apply pythagoras theorem to this triangle and obtain the value of x. (x^2) + (x+30)^2 = (x+60)^2. By solving this equations and cancelling some terms we are left with a simple quadratic equation. Which is x^2 -60x -2700=0. This can be factorized be splitting the mid term and the value of x turns out to be 90. Hence shortest side =90.
sachin
6 years ago
By prepainga diagram, sides x, x+30 and x+60 form a right angle triange inside the rectangle. We can apply pythagoras theorem to this triangle and obtain the value of x. (x^2) + (x+30)^2 = (x+60)^2. By solving this equations and cancelling some terms we are left with a simple quadratic equation. Which is x^2 -60x -2700=0. This can be factorized be splitting the mid term and the value of x turns out to be 90